How much heat is required to change a 47.4 g ice cube from ice at -11.8°C to water at 68°C? (if necessary, use cice=2090 J/kg°C and csteam= 2010 J/kg°C)

How much heat is required to change a 47.4 g ice cube from ice at -11.8°C to steam at 115°C?
Answer #1

specific heat of ice lice = 2090 J/kg°C = 2.0903/9°C specific heat of steam Csteam = 2010 JIK70€ = 2.010 J[°C -11.8°cice to o115° steam is 8 = itezte 4 +95 +86 0° water to 100°c water heat required Bq = m Gwater (Tf - Ti) Tf-Ti = 100-0 = 100°. → 84